Find the equation of a line that is parallel to 3x + 4y = 5 and containing the point (5, 0). Please show all of your work and write your answer in slope-intercept form
Answer by mananth(16946)

3 x + 4 y = 5
Find the slope of this line
make y the subject
4 y = -3 x + 5
Divide by 4
y = - 3/4 x + 1.25
Compare this equation with y=mx+b
slope m = - 3/4
The slope of a line parallel to the above line will be the same
The slope of the required line will be - 3/4
m= - 3/4 ,point ( 5 , 0 )
Find b by plugging the values of m & the point in
y=mx+b
0 = -15/ 4 + b
b= 15/4
m= - 3/4
Plug value of the slope and b in y = mx +b
The required equation is y = - 3/4 x + 15/ 4

m.ananth@hotmail.ca